OpenStudy (anonymous):

pls sove this by partial fraction method(x^2+1)/x(x-1)(x+1)

OpenStudy (anonymous):

Start by expanding (x-1)(x+1) to get x^2-1. Then try this anzats (or estimation): A/x + (Bx+c)/(x^2-1). You can then cross multiply to get: \[x ^{2}+1=A(x^{2}-1)+x(Bx+C)\] Is that enough to work with? I don't know what you mean by "solve" in this context but if you try different x values in the above - or expand out the LHS - you can find A, B & C, which you just sub back in the anzats to find the partial fractions. I hope that helps :)

OpenStudy (anonymous):

Here A= -1, B=2, C=0

OpenStudy (anonymous):

i have tried and got the\[(1\div x)+1\div(x+1)+1\div(x-1)\] answer ...is this right

OpenStudy (anonymous):

i got a=1 b=1 and c=1

OpenStudy (anonymous):

You may be right but I found A by x=0 and B & C by considering x = 1 and x = -1.

OpenStudy (anonymous):

and my text says 1/(x-1)+1/(x+1)-1/x

OpenStudy (anonymous):

We are both correct.

OpenStudy (anonymous):

o.k will try it once again ....thanks alot

OpenStudy (anonymous):

i think A=-1,,B=C=1

OpenStudy (anonymous):

so the text book says A=-1 ok

OpenStudy (anonymous):

i mean A=1... ok

OpenStudy (anonymous):

no A= -1

OpenStudy (anonymous):

B=C=1

OpenStudy (anonymous):

yeah A=-1

OpenStudy (anonymous):

ok as long as you got the correct solution and answer....LOL

OpenStudy (anonymous):

When x=0, 1=A(-1) <=> A=-1. When x=1, 2=A(1-1) +(B+C) and when x=-1, 2=A(1-1) +(B-C) <=> C=0 & B=2.

OpenStudy (anonymous):

mhhhmm thanks both of you

OpenStudy (anonymous):

but 0=Bx - Cx therefore 0=x(B-C) B=C ..If B=1=C

OpenStudy (anonymous):

Where did you get 0=Bx-Cx from?

OpenStudy (anonymous):

expand x^2 + 1 = A(x-1)(x+1) + Bx(X+1) +Cx(x-1)

OpenStudy (anonymous):

That is incorrect: C is the constant term. There is no Cx.

OpenStudy (anonymous):

=Ax^2-A + Bx^2+Bx+Cx^2-Cx

OpenStudy (anonymous):

mark the quation you have given might be correct ....can please elaborate for the value of A

OpenStudy (anonymous):

WHAT VALUE OF "X" DID YOU TAKE TO FIND THE VALUE OF "A"

OpenStudy (anonymous):

No; (x+1)/(x^2-1) - 1/x does not work. But (2x)/(x^2-1) - 1/x does.

OpenStudy (anonymous):

from (x^2 +1) A B C ------- = --- + --- + ---- x(x-1)(x+1) x x-1 x+1

OpenStudy (anonymous):

I was not using that anzats lol. rahulyadav was correct.

OpenStudy (anonymous):

x^2 + 1 = A(x-1)(x+1) + Bx(x+1) + Cx(x-1)

OpenStudy (anonymous):

Here A=-1 and B=C=1.

OpenStudy (anonymous):

mark o. to find value of A what value of x did you take???

OpenStudy (anonymous):

Try x=0.

OpenStudy (anonymous):

When x=0, 1=A(-1)(1) <=> A=-1.

OpenStudy (anonymous):

i am not confident about the value of which i got cause the coefficient of A ALWAYS COMES DOWN TO "0" IF I PUT X=1 OR X=-1

OpenStudy (anonymous):

=A(x^2 -1)+B(x^2 +x) +C(x^2-x) x^2+1 = Ax^2 - A + Bx^2 + Bx + Cx^2 -Cx 1=-A......A=-1

OpenStudy (anonymous):

That does not matter if you have already found A using x=0.

OpenStudy (anonymous):

0=Bx-Cx 0=x(B-C) 0=(B-C)... B=C

OpenStudy (anonymous):

so finally both the values work for me but which value should i use?...x=1,x=-1,or x=0????

OpenStudy (anonymous):

All three.

OpenStudy (anonymous):

yes it works also

OpenStudy (anonymous):

The purpose of using x values is to discover what A, B & C are. The anzats mark put up is called an identity. It works for all x values.

OpenStudy (anonymous):

thanks both of you for giving so much time to this problem

OpenStudy (anonymous):

i just did it a long way.... and you can do it a short cut method like what shaun did....

OpenStudy (anonymous):

ok welcome and good luck

OpenStudy (anonymous):

Strictly speaking, the "=" sign should have three lines to make that clear. When writing an anzats for partial fractions, try using the three lines to remind you that it holds for any x. And you're welcome :)

OpenStudy (anonymous):

yes correct you can use 3 lines in equal sign